# Privacy-Preserved Neural Graph Databases

### v4

@@ Abstract.
In the era of large language models (LLMs), efficient and accurate data retrieval has become increasingly crucial for the use of domain-specific or private data in the retrieval augmented generation (RAG). Neural graph databases (NGDBs) have emerged as a powerful paradigm that combines the strengths of graph databases (GDBs) and neural networks to enable efficient storage, retrieval, and analysis of graph-structured data which can be adaptively trained with LLMs. The usage of neural embedding storage and Complex neural logical Query Answering (CQA) provides NGDBs with generalization ability. When the graph is incomplete, by extracting latent patterns and representations, neural graph databases can fill gaps in the graph structure, revealing hidden relationships and enabling accurate query answering. Nevertheless, this capability comes with inherent trade-offs, as it introduces additional privacy risks to the domain-specific or private databases. Malicious attackers can infer more sensitive information in the database using well-designed queries such as from the answer sets of where Turing Award winners born before 1950 and after 1940 lived, the living places of Turing Award winner Hinton are probably exposed, although the living places may have been deleted in the training stage due to the privacy concerns. In this work, we propose a privacy-preserved neural graph database (P-NGDB) framework to alleviate the risks of privacy leakage in NGDBs. We introduce adversarial training techniques in the training stage to enforce the NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring sensitive information through combinations of multiple innocuous queries. Extensive experimental results on three datasets show that our framework can effectively protect private information in the graph database while delivering high-quality public answers responses to queries.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Privacy protectionsccs: Information systems Data mining
@@ 1. Introduction
Graph DataBases (GDBs) play a crucial role in storing, organizing, and retrieving structured relational data and support many traditional data-intensive applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. In the era of Large Language Models (LLMs), the role of GDBs gets increasingly important because of the Retrieval Augmented Generation (RAG) paradigm, where the LLM agents can be significantly enhanced by external GDBs such as Knowledge Graphs (KGs). [^1] https://openai.com/research/emergent-tool-use [^2] https://docs.llamaindex.ai/en/stable/examples/agent/openai_agent.html [^3] https://www.langchain.com/use-case/agents Such interaction of LLMs and GDBs enables new possibilities by creating interactive natural language interfaces over structural data for domain-specific applications.
@@
To address the challenges, extending the concept of graph DBs, Ren et al. (Ren et al., 2023) propose neural graph databases (NGDBs) which combine the flexibility of graph data models with the computational capabilities of neural networks, enabling effective and efficient representation, storage, and analysis of interconnected data. As shown in Figure 1, NGDBs are designed with two modules: neural graph storage providing unified storage for diverse entries in an embedding space and neural query engine searching answers to input complex queries from the unified storage (Ren et al., 2023). By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA) (Arakelyan et al., 2021), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018). Query encoding methods are commonly used in CQA. These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query (or set) embeddings and answer embeddings. For example, GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020), and Q2P (Bai et al., 2022) encode queries as vectors, a hyper-rectangle, and a set of particles, respectively. Ideally, CQA can be jointly trained with RAG so that the language model agents can better adapt to domain-specific data (Lee et al., 2019; Guu et al., 2020).
@@
While NGDBs and their CQA ability have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs. One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge and enriches the retrieved information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. Figure 1 demonstrates an example of privacy attacks via malicious complex queries: the living place information of an individual. In this case, Hinton as an example, is considered to be private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place but conducting well-designed related queries. The answers denoted in red may leak sensitive information by accident.
@@
Graph privacy has consistently been a significant concern in the field of graph research. Multiple studies demonstrate the vulnerability of graph neural networks to various privacy leakage issues. For instance, graph embeddings are vulnerable to attribute inference attacks (Duddu et al., 2020; Gong and Liu, 2018; Zhang et al., 2022) and link prediction attacks (He et al., 2021b; Wu et al., 2022), and various protection methods are proposed (Hu and Song, 2023; Hu et al., 2022; Wang et al., 2021a). However, these works focus on graph representation learning and have the assumption that the graph embeddings can be reached by the attackers which are not typical scenarios in NGDBs. NGDBs often provide CQA service and do not publish the learned embeddings, and attackers can only infer private information using thoughtfully designed queries. Furthermore, these works solely study node classification and link prediction problems, leaving the query answering privacy in NGDBs unexplored.
@@ 2.2. Graph Privacy
Various privacy protection methods have been proposed to preserve privacy in graph representation models. Anonymization techniques (Liu and Terzi, 2008; Rossi et al., 2015; Hoang et al., 2020; Zheleva and Getoor, 2007) are applied in graphs to reduce the probability of individual and link privacy leakage. Graph summarization (Hay et al., 2008) aims to publish a set of anonymous graphs with private information removed. Learning-based methods are proposed with the development of graph representation learning, (Wang et al., 2021a; Liao et al., 2020; Li et al., 2020) regard graph representation learning as the combination of two sub-tasks: primary objective learning and privacy preservation, and use adversarial training to remove the sensitive information while maintaining performance in the original tasks. Meanwhile, some methods (Liu et al., 2022; Oh et al., 2022; Hu and Song, 2023) disentangle the sensitive information from the primary learning objectives. Additionally, differential privacy (Kasiviswanathan et al., 2013; Day et al., 2016; Shen and Yu, 2013) introduces noise into representation models and provides privacy guarantees for the individual privacy of the datasets (Daigavane et al., 2021). Though noise introduced by differential privacy protects models from privacy leakage, it can also impact the performance of the original tasks significantly. Federated learning utilizes differential privacy and encryption to prevent the transmission of participants’ raw data and is widely applied in distributed graph representation learning (Peng et al., 2021; He et al., 2021a). However, federated learning cannot be applied to protect intrinsic private information in the representation models.
@@ 3.1. Preliminary
We denote a knowledge graph as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A})$, where $\mathcal{V}$ denotes the set of vertices representing entities and attribute values in the knowledge graph. $\mathcal{R}$ denotes the set of relations in the knowledge graph. $\mathcal{A}$ denotes the set of attributes that can be split into private attributes and public attributes: $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a\in\mathcal{A}$ is a binary function defined as $a:\mathcal{V}\times\mathcal{V}\rightarrow\{0,1\}$ describing the attribute of entities, where $a(u,x)=1$ denotes the entity $u$ has the attribute $a$ of value $x\in\mathcal{V}$ and $a(u,x)=0$ otherwise. To prevent privacy leakage, the private attribute of entities, we denote the $\mathcal{A}_{\text{private}}$ cannot be exposed and should be protected from inferences.
@@ 3.2. Problem Formulation
Suppose that there is a graph $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A})$, where part of entities’ attributes are regarded as private information $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a_{p}\in\mathcal{A}_{\text{private}}$ denotes the entity $u$ has the sensitive attribute $a_{p}$ with value $x$ if $a_{p}(u,x)=1$. Due to the variation in privacy requirements, an attribute can be handled differently for various entities. NGDBs store the graph in embedding space and can be queried with arbitrary complex logical queries. Due to the generalization ability of NGDBs, attackers can easily infer sensitive attributes $\mathcal{A}_{\text{private}}$ utilizing complex queries. To protect private information, NGDBs should retrieve obfuscated answers when queried with privacy risks while preserving high accuracy in querying public information.
@@ 4. Innocuous yet Private Query
Figure 2. An example of query demonstrating the retrieved privacy-threatening query answers. The red node denotes privacy risks. (A) The logic knowledge graph query involves privacy information. (B) An example of a complex query in the knowledge graph. Toronto is regarded as a privacy-threatening answer as it has to be inferred by sensitive information.
Figure 3. Example of privacy-threatening answer sets computation in projection, intersection, and union. Green nodes denote non-private answers, orange nodes denote privacy-threatening answers, and orange-green nodes denote different privacy risks in subsets. Red dashed arrows denote privacy projection. The range demarcated by the red dashed lines denotes privacy-threatening answer sets. The answers circled in red dashed line are at risk to leak privacy.
@@ 4.1. Computational Graph
  • Projection: Projection has various types for attribute- and value-sensitive queries (Bai et al., 2023a). Given a set of entities $S$, a relation between entities $r\in\mathcal{R}$, an attribute type $a$, a set of values $S\subset\mathcal{N}$, projection describes the entities or values can be achieved under specified relation and attribute types. For example, attribute projection is denoted as $P_{a}(S)=\{x\in\mathcal{N}|v\in S,a(v,x)=1\}$ describing the values that can be achieved under the attribute type $a$ by any of the entities in $S$. Furthermore, we have relational projection $P_{r}(S)$, reverse attribute projection $P_{a}^{-1}(S)$ and numerical projection $P_{f}(S)$. We denote all these projections as $P(S)$;
@@ 4.2. Privacy Threatening Query Answers
Projection: A projection operator’s output will be regarded as private if the operation is applied to infer private attributes. Assume that $M=P(S)$ and input $S=S_{\text{private}}\cup S_{\text{public}}$. $P(S)=P(S_{\text{private}})\cup P(S_{\text{public}})$. Without loss of generalization, we discuss the formal definition of projection private answer set $M_{\text{private}}$ in two scenarios. For the projection on public inputs:
@@
$P(u,m)\neq P(v,m),\forall u\in S/\{v\}\}.$
@@
The answers of projection on private inputs are privacy threatening answers as they can expose the existence of private inputs. As shown in Figure 3(A), all the red nodes within the answer sets can be solely deduced from private attribute links and are classified as query answers posing privacy risks. On the other hand, nodes such as the green-red node, while accessible through private projection, can also be inferred from public components, thereby carrying lower privacy risks.
Intersection: An intersection operator’s output will be regarded as private if the answer belongs to any of the private answer sets. Given a set of answer sets $M^{1},M^{2},...,M^{n}$, where each answer set $M^{i}=M^{i}_{\text{private}}\cup M^{i}_{\text{public}}$, after intersection operator, $M_{\text{private}}=\cap_{i=1}^{n}M^{i}-\cap_{i=1}^{n}M^{i}_{\text{public}}$. As shown in Figure 3(B), the green-red node is categorized as a privacy-threatening answer because it denotes the existence in the answer subsets.
Union: A union operator’s output will be regarded as private if the answer is the element that belongs to the private answer set of computational subgraphs while not belonging to public answer sets. Given a set of answer sets $M^{1},M^{2},...,M^{n}$, $M_{\text{private}}=\cup_{i=1}^{n}M^{i}_{\text{private}}-\cup_{i=1}^{n}M^{i}_{\text{public}}$. As shown in Figure 3(C), all the red nodes within the answer sets can only be inferred from private attribute links and are classified as privacy-threatening answers.
@@ 5.2. Learning Objective
(6) $L_{u}=-\frac{1}{N}\sum_{v\in\mathcal{N}}\log p(q,v),$
where $\mathcal{N}$ is the public answer set for the given query.

### v5

@@ Abstract.
In the era of large language models (LLMs), efficient and accurate data retrieval has become increasingly crucial for the use of domain-specific or private data in the retrieval augmented generation (RAG). Neural graph databases (NGDBs) have emerged as a powerful paradigm that combines the strengths of graph databases (GDBs) and neural networks to enable efficient storage, retrieval, and analysis of graph-structured data which can be adaptively trained with LLMs. The usage of neural embedding storage and Complex neural logical Query Answering (CQA) provides NGDBs with generalization ability. When the graph is incomplete, by extracting latent patterns and representations, neural graph databases can fill gaps in the graph structure, revealing hidden relationships and enabling accurate query answering. Nevertheless, this capability comes with inherent trade-offs, as it introduces additional privacy risks to the domain-specific or private databases. Malicious attackers can infer more sensitive information in the database using well-designed queries such as from the answer sets of where Turing Award winners born before 1950 and after 1940 lived, the living places of Turing Award winner Hinton are probably exposed, although the living places may have been deleted in the training stage due to the privacy concerns. In this work, we propose a privacy-preserved neural graph database (P-NGDB) framework to alleviate the risks of privacy leakage in NGDBs. We introduce adversarial training techniques in the training stage to enforce the NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring sensitive information through combinations of multiple innocuous queries. Extensive experimental results on three datasets show that our framework can effectively protect private information in the graph database while delivering high-quality public answers responses to queries. The code is available at https://github.com/HKUST-KnowComp/PrivateNGDB.
@@
copyright: acmcopyrightjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671678isbn: 979-8-4007-0490-1/24/08ccs: Security and privacy Privacy protectionsccs: Information systems Data mining
@@ 1. Introduction
Graph DataBases (GDBs) play a crucial role in storing, organizing, and retrieving structured relational data and support many traditional data-intensive applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. In the era of Large Language Models (LLMs), the role of GDBs gets increasingly important because of the Retrieval Augmented Generation (RAG) paradigm, [^1] https://openai.com/research/emergent-tool-use [^2] https://docs.llamaindex.ai/en/stable/examples/agent/openai_agent.html [^3] https://www.langchain.com/use-case/agents where the LLM agents can be significantly enhanced by external GDBs such as Knowledge Graphs (KGs) (Edge et al., 2024; Mavromatis and Karypis, 2024; Gao et al., 2023). Such interaction of LLMs and GDBs enables new possibilities by creating interactive natural language interfaces over structural data for domain-specific applications.
@@
NGDBs have been proposed to address these challenges by extending the concept of graph DBs (Ren et al., 2023; Besta et al., 2022). They combine the flexibility of graph data models with the computational capabilities of neural networks, enabling effective and efficient representation, storage, and analysis of interconnected data. As shown in Figure 1, NGDBs are designed with two modules: neural graph storage providing unified storage for diverse entries in an embedding space and neural query engine searching answers to input complex queries from the unified storage (Ren et al., 2023). By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA) (Arakelyan et al., 2021), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018). Query encoding methods are commonly used in CQA. These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query (or set) embeddings and answer embeddings. For example, GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020), and Q2P (Bai et al., 2022) encode queries as vectors, a hyper-rectangle, and a set of particles, respectively. Ideally, CQA can be jointly trained with RAG so that the language model agents can better adapt to domain-specific data (Lee et al., 2019; Guu et al., 2020).
@@
While NGDBs and their CQA ability have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs (Zeighami and Shahabi, 2023; Zeighami et al., 2023). One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge and enriches the retrieved information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. Figure 1 demonstrates an example of privacy attacks via malicious complex queries: the living place information of an individual. In this case, Hinton as an example, is considered to be private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place but conducting well-designed related queries. The answers denoted in red may leak sensitive information by accident.
@@
Graph privacy has consistently been a significant concern in the field of graph research. Multiple studies demonstrate the vulnerability of graph neural networks to various privacy leakage issues. For instance, graph embeddings are vulnerable to attribute inference attacks (Duddu et al., 2020; Gong and Liu, 2018; Zhang et al., 2022) and link prediction attacks (He et al., 2021b; Wu et al., 2022), and various protection methods are proposed (Hu and Song, 2023a; Hu et al., 2022; Wang et al., 2021a). However, these works focus on graph representation learning and have the assumption that the graph embeddings can be reached by the attackers which are not typical scenarios in NGDBs. NGDBs often provide CQA service and do not publish the learned embeddings, and attackers can only infer private information using thoughtfully designed queries. Furthermore, these works solely study node classification and link prediction problems, leaving the query answering privacy in NGDBs unexplored, resulting in the potential privacy leakage problem in LLMs (Li et al., 2023a).
@@ 2.2. Graph Privacy
Various privacy protection methods have been proposed to preserve privacy in graph representation models. Anonymization techniques (Liu and Terzi, 2008; Rossi et al., 2015; Hoang et al., 2020; Zheleva and Getoor, 2007) are applied in graphs to reduce the probability of individual and link privacy leakage. Graph summarization (Hay et al., 2008) aims to publish a set of anonymous graphs with private information removed. Learning-based methods are proposed with the development of graph representation learning, (Wang et al., 2021a; Liao et al., 2020; Li et al., 2020) regard graph representation learning as the combination of two sub-tasks: primary objective learning and privacy preservation, and use adversarial training to remove the sensitive information while maintaining performance in the original tasks. Meanwhile, some methods (Liu et al., 2022; Oh et al., 2022; Hu and Song, 2023a) disentangle the sensitive information from the primary learning objectives. Additionally, differential privacy (Kasiviswanathan et al., 2013; Day et al., 2016; Shen and Yu, 2013) introduces noise into representation models and provides privacy guarantees for the individual privacy of the datasets (Daigavane et al., 2021). Though noise introduced by differential privacy protects models from privacy leakage, it can also impact the performance of the original tasks significantly. To empirically evaluate the privacy leakage problem in representations, some works are proposed to construct benchmarks (Li et al., 2023b). Federated learning utilizes differential privacy and encryption to prevent the transmission of participants’ raw data and is widely applied in distributed graph representation learning (Peng et al., 2021; He et al., 2021a; Hu and Song, 2023b). However, federated learning cannot be applied to protect intrinsic private information in the representation models.
@@ 3.1. Preliminary
We denote a knowledge graph as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A})$, where $\mathcal{V}$ denotes the set of vertices representing entities and attribute values in the knowledge graph. $\mathcal{R}$ denotes the set of relations in the knowledge graph. $\mathcal{A}$ denotes the set of attributes that can be split into private attributes and public attributes: $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a(u,x)\in\mathcal{A}$ is a triplet where $u,x\in\mathcal{V}$, describing the attribute of entities. $a(u,x)$ denotes that the entity $u$ has the attribute $a$ of value $x\in\mathcal{V}$. To prevent privacy leakage, the private attribute of entities, we denote the $\mathcal{A}_{\text{private}}$ cannot be exposed and should be protected from inferences.
@@ 3.2. Problem Formulation
Suppose that there is a graph $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A})$, where part of entities’ attributes are regarded as private information $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a_{p}(u,x)\in\mathcal{A}_{\text{private}}$ denotes that the entity $u$ has the sensitive attribute $a_{p}$ with value $x$. Due to the variation in privacy requirements, an attribute can be handled differently for various entities. NGDBs store the graph in embedding space and can be queried with arbitrary complex logical queries. Due to the generalization ability of NGDBs, attackers can easily infer sensitive attributes $\mathcal{A}_{\text{private}}$ utilizing complex queries. To protect private information, NGDBs should retrieve obfuscated answers when queried with privacy risks while preserving high accuracy in querying public information.
@@ 4. Innocuous yet Private Query
Figure 2. An example of query demonstrating the retrieved privacy-threatening query answers. The orange node denotes privacy risks. (A) The logic knowledge graph query involves privacy information. (B) An example of a complex query in the knowledge graph. Toronto is regarded as a privacy-threatening answer as it has to be inferred by sensitive information.
Figure 3. Example of privacy-threatening answer sets computation in projection, intersection, and union. Green nodes denote non-private answers, orange nodes denote privacy-threatening answers, and green-orange nodes denote different privacy risks in subsets. Red dashed arrows denote privacy projection. The range demarcated by the red dashed lines denotes privacy-threatening answer sets. The answers circled in red dashed line are at risk of leaking privacy.
@@ 4.1. Computational Graph
  • Projection: Projection has various types for attribute- and value-sensitive queries (Bai et al., 2023a). Given a set of entities $S$, a relation between entities $r\in\mathcal{R}$, an attribute type $a$, a set of values $S\subset\mathcal{V}$, projection describes the entities or values can be achieved under specified relation and attribute types. For example, attribute projection is denoted as $P_{a}(S)=\{x\in\mathcal{V}|v\in S,a(v,x)\in\mathcal{A}\}$ describing the values that can be achieved under the attribute type $a$ by any of the entities in $S$; Furthermore, we have relational projection $P_{r}(S)$, reverse attribute projection $P_{a}^{-1}(S)$. We denote all these projections as $P(S)$;
@@ 4.2. Privacy Threatening Query Answers
Projection: A projection operator’s output will be regarded as private if the operation is applied to infer private attributes. Assume that $M=P(S)$ and input $S=S_{\text{private}}\cup S_{\text{public}}$. $P(S)=P(S_{\text{private}})\cup P(S_{\text{public}})$, we use $p$ denote the corresponding attribute of projection $P$. Without loss of generalization, we discuss the formal definition of projection private answer set $M_{\text{private}}$ in two scenarios. For the projection on public inputs:
@@
$p(u,m)\neq p(v,m),\forall u\in S/\{v\},p(u,m)\notin\mathcal{A}_{\text{private}}\}.$
@@
The answers of projection on private inputs are privacy threatening answers as they can expose the existence of private inputs. As shown in Figure 3(A), all the orange nodes within the answer sets can be solely deduced from private attribute links and are classified as query answers posing privacy risks. On the other hand, nodes such as the green-orange node, while accessible through private projection, can also be inferred from public components, thereby carrying lower privacy risks.
Intersection: An intersection operator’s output will be regarded as private if the answer belongs to any of the private answer sets. Given a set of answer sets $M^{1},M^{2},...,M^{n}$, where each answer set $M^{i}=M^{i}_{\text{private}}\cup M^{i}_{\text{public}}$, after intersection operator, $M_{\text{private}}=\cap_{i=1}^{n}M^{i}-\cap_{i=1}^{n}M^{i}_{\text{public}}$. As shown in Figure 3(B), the green-orange node is categorized as a privacy-threatening answer because it denotes the existence in the answer subsets.
Union: A union operator’s output will be regarded as private if the answer is the element that belongs to the private answer set of computational subgraphs while not belonging to public answer sets. Given a set of answer sets $M^{1},M^{2},...,M^{n}$, $M_{\text{private}}=\cup_{i=1}^{n}M^{i}_{\text{private}}-\cup_{i=1}^{n}M^{i}_{\text{public}}$. As shown in Figure 3(C), all the orange nodes within the answer sets can only be inferred from private attribute links and are classified as privacy-threatening answers.
@@ 5.2. Learning Objective
(6) $L_{u}=-\frac{1}{N}\sum_{v\in\mathcal{M}_{\text{public}}^{q}}\log p(q,v),$
where $\mathcal{M}_{\text{public}}^{q}$ is the public answer set for the given query.